OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2 .

OpenStudy (miguel980):

\[-4=1/2(1) + b\]\[-4=1/2 + b\]\[-1/2 =-1/2\]\[-4 \frac{ 1 }{ 2 }= b\]

OpenStudy (anonymous):

y +4 = (1/2) (x-1)